# Dynamics and fluid-structure interaction in turbulent flows within and above flexible canopies

### 21 Dec 2023 v1

@@ Abstract
Flexible canopy flows are often encountered in natural scenarios, eg., when crops sway in the wind or when submerged kelp forests are agitated by a marine current. Here, we provide a detailed characterisation of the turbulent flow developed above and between the flexible filaments of a fully submerged dense canopy and we describe their dynamical response to the turbulent forcing. We investigate a wide range of flexibilities, encompassing the case in which the filaments are completely rigid and standing upright as well as that where they are fully compliant to the flow and deflected in the streamwise direction. We are thus able to isolate the effect of the canopy flexibility on the drag and on the inner-outer flow interactions, as well as the two flapping regimes of the filaments already identified for a single fiber. Furthermore, we offer a detailed description of the Reynolds stresses throughout the wall-normal direction resorting to the Lumley triangle formalism, and we show the multi-layer nature of turbulence inside and above the canopy. The relevance of our investigation is thus twofold: the fundamental physical understanding developed here paves the way towards the investigation of more complex and realistic scenarios, while we also provide a thorough characterisation of the turbulent state that can prove useful in the development of accurate turbulence models for RANS and LES.
@@ 1 Introduction
The chaotic motion of air in and above plant canopies has been systematically investigated and modelled from the second half of the 20th century. In their seminal review, Raupach & Thom (1981) laid solid foundations for the study of turbulence and transport in canopy flows, providing a first characterisation of those phenomena and reviewing the most successful approaches for the prediction of the mean flow and turbulence intensity. An exhaustive description of the mean flow features and a detailed characterisation of the key turbulent quantities was later offered by Finnigan (2000). Finnigan (2000) also introduced a phenomenological model for the sustainment of turbulence, reliant on the inviscid instability of the shear layer generated by the drag discontinuity at the tip of a dense submerged canopy. The relevance of such shear layer was first observed by Raupach et al. (1996), while its peculiar nature was successively highlighted by Ghisalberti & Nepf (2004), who noted that it does not grow continuously downstream as a free shear layer, but rather reaches a finite thickness set by the rate of momentum exchange between the flow and the solid structures. Notwithstanding this noticeable difference, it is yet passible of a Kelvin-Helmholtz like instability which induces the formation of elongated spanwise vortices (“rollers”) controlling the exchange of mass and momentum between the canopy and the outer flow (Nepf, 2012a; Chowdhuri et al., 2022). The further instability of those rollers is held responsible for the formation of secondary vortices aligned with the flow, causing intense sweeps often observed to penetrate the canopy. Alongside the theoretical approach, there has been a flourishing of models to predict various quantities of interest based on experimental measurements. In particular, the canopy drag coefficient is relevant in most engineering applications and exhibits a direct dependence from the Reynolds and the Froude numbers, as demonstrated by Liu & Zeng (2016), Mossa et al. (2021) and Rubol et al. (2018), who further investigated its dependence from the canopy permeability. Additional models for other flow quantities are summarised by Brunet (2020), who offers a physical overview and presents a historical summary of the evolution of the field. Recently, Vieira et al. (2023) have introduced a simplified model of the unidirectional flow over a canopy capable of accounting for the shear layer instability above its tip, while Conde-Frias et al. (2023) have developed an experimentally-validated approach to predict the boundary layer thickness at the seafloor based upon its negative correlation with the turbulent kinetic energy within the layer.
@@
The picture becomes more complex in the case of flexible slender elements, as their structural dynamics needs to be accounted for, and a significant portion of the turbulent kinetic energy is generated by the “waving contribution” originated from the correlation between the hydrodynamic drag and the waving motion of the stems (He et al., 2022). Jin et al. (2016) laid the foundations for understanding the dynamical response of flexible and slender elements forced by a turbulent flow. On top of that, Rosti et al. (2018a) developed a phenomenological theory to describe the dynamics of free flexible fibres in homogeneous isotropic turbulence and validated it numerically. They identified two regimes of motion: one in which the fibres are slaved to the turbulent fluctuations of the flow and one in which they exhibit their natural response; Olivieri et al. (2021) later supported the result with a wider set of numerical simulations. We confirmed the existence of those two regimes also in the case of a clamped flexible fibre in wall turbulence (Foggi Rota et al., 2023), nevertheless noticing a significant difference in the dominant oscillation frequency for the turbulence dominated regime. In this case, in facts, the flapping state of the fiber non-trivially relates to the largest scale of the flow and not to the turbulent eddies of comparable size, as found by Olivieri et al. (2021). In the case of a flexible canopy, the individual dynamics of the elements is altered (Fu et al., 2023) and a collective dynamics (honami/monami) emerges on top of that, as measured experimentally by Py et al. (2006) in the case of a crop field driven by the wind. The state-of-the-art direct numerical simulations (DNS) performed by Monti et al. (2023) shed light on the topic, highlighting how such collective motion is decoupled from the structural natural response of the filaments and independent from their rigidity, but instead it is driven only by the turbulent fluctuations of the flow. Such causal dependence is coherent with the previous results of the large eddy simulations from Tschisgale et al. (2021) and Wang et al. (2022a).
@@
In this work we investigate the turbulent motion of the fluid and the dynamics of the flexible filaments constituting a dense canopy. We build on top of an extended version of the DNS database generated by Monti et al. (2023), where each filament is modelled individually by means of an immersed boundary method (first introduced by Goldstein et al., 1993) in the Lagrangian formulation described by Yu (2005) and Huang et al. (2007). Consistently with such approach, the filaments are therefore modelled as inextensible beams. We provide a detailed characterisation of the flow above and within the canopy, as well as describing the interaction between the two. We also assess the consequences on the flow of a variation in the filament density and analyse the effects induced by their motion. From a structural standpoint, we characterise the individual motion of the filaments in the canopy and reconcile it with the picture we previously described for single isolated filaments (Foggi Rota et al., 2023). The remaining part of this paper is organised as follows: in §2 we describe the setup of our simulations and the numerical methods employed. The dynamics of the flow is characterised in §3 where, after reporting relevant mean quantities (§3.1), we investigate the energy spectra (§3.2) and the Lumley triangle (§3.3) throughout the whole domain. We further explore the interaction between the inner and the outer flows by carrying out a quadrant analysis (§3.4) at the canopy tip. The consequences on the flow of a variation in the density of the filaments are also assessed in §3.5 and the effects of their flapping motion are explored by “freezing” them in their instantaneous deflected configuration in §3.6. In §4 we characterise the structural dynamics of the filaments by analysing their individual motion (§4.1) and describing their flapping state (§4.2). Finally, in §5, we summarise the main outcomes of our investigation and critically discuss their implications, with few remarks on future developments.
@@ 2 Setup and methods
Our simulations are carried out in an open channel described by means of a right handed Cartesian reference frame with the $x$ axis oriented along the streamwise direction and the $y$ axis perpendicular to the bottom wall. The computational domain is therefore a box of volume $L_{x}\times L_{y}\times L_{z}=2\pi H\times H\times 1.5\pi H$, where we enforce the no-slip and no-penetration boundary conditions at the bottom face; no-penetration and stress-free conditions are instead imposed at the top face in the same fashion of Calmet & Magnaudet (2003), while the wall-parallel directions are treated as periodic. The stems of the flexible filaments constituting the canopy are vertically clamped to the bottom wall and protrude upward into the flow as in figure 1.
To discretise the fluid flow we adopt an Eulerian grid made of $N_{x}\times N_{y}\times N_{z}=1152\times 384\times 864$ points homogeneously distributed along the periodic directions, while a non homogeneous distribution is adopted along the $y$ axis in order to correctly capture the sharp variation of the velocity at the canopy tip. In particular, as shown in figure 2, we employ a finer and locally uniform resolution in the region containing the canopy, smoothly transitioning to a wider wall-normal spacing above that.
@@
where two forcing terms, $\mathbf{f}_{\mathbf{\rm{fib}}}$ and $\mathbf{f}_{\mathbf{\rm{for}}}$, are introduced to account for the fluid-structure interaction and to sustain a fully developed turbulent flow at constant Reynolds number $Re_{b}=U_{b}H/\nu=5000$, with $U_{b}$ the average flow velocity in the domain. The balance equations, along with the set of boundary conditions highlighted above, constitute a well posed problem that we tackle numerically by means of our in-house solver, Fujin (https://groups.oist.jp/cffu/code). We adopt second order central finite differences to discretise the velocity and the pressure on a staggered Cartesian grid, resorting to a second order Adams-Bashforth scheme for time stepping within a projection-correction approach (Kim & Moin, 1985). The Poisson equation is efficiently solved with a Fast Fourier Transform (FFT) based algorithm (Dorr, 1970) and the whole code is parallelised using the Message Passing Interface (MPI) and the 2decomp library.
Figure 2: Wall-normal Eulerian grid adopted throughout our simulations. We show (panel $a$) the incremental number of grid points $n_{y}$ and (panel $b$) the wall-normal grid spacing $\Delta y$ against the wall distance.
The canopy is constituted by $n_{x}\times n_{z}=144\times 108$ filaments of length $h=0.25H$ and diameter $d\approx 5.45\cdot 10^{-3}H$, placed in a semi-random arrangement to avoid preferential flow channeling effects. The bottom wall is therefore divided in regular tiles of area $\Delta S^{2}=(L_{x}/n_{x})\times(L_{z}/n_{z})$ and a filament is placed at a random position within each one. We thus attain a nominal solidity value of $\lambda=hd/\Delta S^{2}\approx 1.43$, laying well within the dense canopy regime (Monti et al., 2020). Different canopies are produced on varying the Cauchy number, $Ca=(\rho_{f}dh^{3}U_{b}^{2})/(2\gamma)$ (representing the ratio between the deforming force exerted by the fluid and the elastic restoring force opposed by the filaments) and the volume density ratio between the filaments and the fluid, $\rho_{s}/\rho_{f}$. Here, $\gamma$ is the bending rigidity of the filaments. Most of our study focuses on six canopies characterised by $\rho_{s}/\rho_{f}=1.0+1.46\cdot 10^{-3}$ spanning $Ca\in\{0,1,10,25,50,100\}$, while for few specific purposes discussed in §3.5 and §4 we also consider a more flexible case with $Ca=500$ and two additional ones with a different density ratio, $\rho_{s}/\rho_{f}\in\{1.0+1.46\cdot 10^{-1},1.0+1.46\cdot 10^{-2}\}$ at $Ca=25$. The filaments are represented as mono-dimensional entities, discretised into a line of Lagrangian points, that obey a generalisation of the Euler-Bernoulli beam model allowing for finite deflections, but retaining the inextensibility constraint. In the rigid canopy case (i.e., $Ca=0$) each filament is made of $n_{L}=81$ Lagrangian points, attaining a spatial resolution $\Delta s=h/(n_{L}-1)$ comparable to the Eulerian grid spacing in the wall-normal direction. In the flexible canopy cases, such spatial resolution would impose a too strict constraint on the time step and we therefore reduce $n_{L}$ to 32: this proves acceptable as the flexibility makes the filaments more compliant to the flow and the velocity difference between the two phases is therefore reduced.
Banaei et al. (2020) used our same approach to model the dynamics of flexible and inextensible filaments. It consists of an extended version of the distributed-Lagrange-multiplier/fictitious-domain (DLM/FD) formulation of the continuum equations introduced by Yu (2005). Denoting with $\mathbf{X}(s,t)$ the position of a point on the neutral axis of a filament as a function of the curvilinear abscissa $s$ and time $t$, and introducing the linear density difference between the filament and the fluid $\Delta\tilde{\rho}=(\rho_{s}-\rho_{f})\pi d^{2}/4$, its structural dynamics is described by
@@
where $T$ is the tension enforcing the inextensibility and $\mathbf{F}$ is the fluid-structure coupling term. We complement equations (3,4) with an appropriate set of boundary conditions, imposing $\mathbf{X}\rvert_{s=0}=\mathbf{X_{0}}$ along with ${\partial\mathbf{X}}/{\partial s}\rvert_{s=0}=(0,1,0)$ at the clamp and ${\partial^{3}\mathbf{X}}/{\partial s^{3}}\rvert_{s=h}={\partial^{2}\mathbf{X}}/{\partial s^{2}}\rvert_{s=h}=\mathbf{0}$ along with $T\rvert_{s=h}=0$ at the free end, and solve them following the approach of Huang et al. (2007). Nevertheless, here, the bending term is treated implicitly as in Banaei et al. (2020) to allow for a larger time step. The set of Lagrangian equations introduced above, in the absence of any external forcing, is passible of a normal mode analysis yielding the natural frequency $f_{nat}\approx 3.516/(dh^{2})\sqrt{\gamma/(\rho_{s}\pi^{3})}$, which plays a significant role in determining the dynamical response of the filaments to the fluid. In principle, filament-to-filament and filament-to-wall collisions need to be accounted for; nevertheless, after extensive testing of elastic and inelastic collision models, filament-to-filament collisions are found to be statistically irrelevant from a dynamical standpoint and are therefore neglected, while filament-to-wall collisions become marginally relevant only for large values of $Ca$. We therefore account for them by means of an inelastic collision model, applying a repulsive force to the filament points approaching the wall within a range of four grid points.
The coupling between the fluid and the structure is attained spreading over the Eulerian grid points the force distribution computed by means of a Lagrangian immersed boundary method (Peskin, 2002), ensuring the no-slip condition $\partial\mathbf{X}/\partial t=\mathbf{u}[\mathbf{X}(s,t),t]$ at the Lagrangian points representing the filaments. The interface between the fluid and the structure is therefore not sharply captured, but diffused between the grid points through the action of a window function. Based on this, we can estimate an effective filament diameter of about four grid points, thus yielding a local Reynolds number $Re_{d}=U_{mc}d/\nu\approx 10$ based on the mean velocity scale $U_{mc}$ within the canopy.
The ability of our numerical setup and methods to correctly describe the dynamics of a whole canopy made of flexible filaments, without introducing spurious effects due to the finite size of the domain or the grid, is extensively assessed in Monti et al. (2023), with references. There, the grid refinement study is provided along with further validation details for a turbulent canopy flow.

### 14 May 2024 v2

@@ Abstract
Flexible canopy flows are often encountered in natural scenarios, e.g., when crops sway in the wind or when submerged kelp forests are agitated by marine currents. Here, we provide a detailed characterisation of the turbulent flow developed above and between the flexible filaments of a fully submerged dense canopy and we describe their dynamical response to the turbulent forcing. We investigate a wide range of flexibilities, encompassing the case in which the filaments are completely rigid and standing upright as well as that where they are fully compliant to the flow and deflected in the streamwise direction. We are thus able to isolate the effect of the canopy flexibility on the drag and on the inner-outer flow interactions, as well as the two flapping regimes of the filaments already identified for a single fiber. Furthermore, we offer a detailed description of the Reynolds stresses throughout the wall-normal direction resorting to the Lumley triangle formalism, and we show the multi-layer nature of turbulence inside and above the canopy. The relevance of our investigation is thus twofold: the fundamental physical understanding developed here paves the way towards the investigation of more complex and realistic scenarios, while we also provide a thorough characterisation of the turbulent state that can prove useful in the development of accurate turbulence models for RANS and LES.
@@ 1 Introduction
The chaotic motion of air in and above plant canopies has been systematically investigated and modelled from the second half of the 20th century. In their seminal review, Raupach & Thom (1981) laid solid foundations for the study of turbulence and transport in canopy flows, providing a first characterisation of those phenomena and reviewing the most successful approaches for the prediction of the mean flow and turbulence intensity. An exhaustive description of the mean flow features and a detailed characterisation of the key turbulent quantities was later offered by Finnigan (2000). Finnigan (2000) also introduced a phenomenological model for the sustainment of turbulence, reliant on the inviscid instability of the shear layer generated by the drag discontinuity at the tip of a dense submerged canopy. The relevance of such shear layer was first observed by Raupach et al. (1996), while its peculiar nature was successively highlighted by Ghisalberti & Nepf (2004), who noted that it does not grow continuously downstream as a free shear layer, but rather reaches a finite thickness set by the rate of momentum exchange between the flow and the solid structures. Notwithstanding this noticeable difference, it is yet passible of a Kelvin-Helmholtz like instability which induces the formation of elongated spanwise vortices (“rollers”) controlling the exchange of mass and momentum between the canopy and the outer flow (Nepf, 2012a; Chowdhuri et al., 2022). The further instability of those rollers is held responsible for the formation of secondary vortices (Finnigan et al., 2009), organised in trains of head-up and head-down hairpins aligned with the flow, causing intense sweeps often observed to penetrate the canopy. The slow fluid drawn from inside the canopy in between the legs of the hairpins, instead, gives rise to elongated regions of low streamwise velocity close to the canopy tip. The rollers, hairpins and velocity streaks generated by the shear layer instability dictate the structure of turbulence throughout the flow, as happens also in highly permeable porous media Manes et al. (2011).
Alongside the theoretical approach, there has been a flourishing of models to predict various quantities of interest based on experimental measurements. In particular, the canopy drag coefficient is relevant in most engineering applications and exhibits a direct dependence from the Reynolds and the Froude numbers, as demonstrated by Liu & Zeng (2016), Mossa et al. (2021) and Rubol et al. (2018), who further investigated its dependence from the canopy permeability. Additional models for other flow quantities are summarised by Brunet (2020), who offers a physical overview and presents a historical summary of the evolution of the field. Recently, Vieira et al. (2023) have introduced a simplified model of the unidirectional flow over a canopy capable of accounting for the shear layer instability above its tip, while Conde-Frias et al. (2023) have developed an experimentally-validated approach to predict the boundary layer thickness at the seafloor based upon its negative correlation with the turbulent kinetic energy within the layer.
@@
The picture becomes more complex in the case of flexible slender elements, as their structural dynamics needs to be accounted for, and a significant portion of the turbulent kinetic energy is generated by the “waving contribution” originated from the correlation between the hydrodynamic drag and the waving motion of the stems (He et al., 2022). Jin et al. (2016) laid the foundations for understanding the dynamical response of flexible and slender elements forced by a turbulent flow. On top of that, Rosti et al. (2018a) developed a phenomenological theory to describe the dynamics of free flexible fibres in homogeneous isotropic turbulence and validated it numerically. They identified two regimes of motion: one in which the fibres are slaved to the turbulent fluctuations of the flow and one in which they exhibit their natural response; Olivieri et al. (2021) later supported the result with a wider set of numerical simulations. We confirmed the existence of those two regimes also in the case of a clamped flexible fibre in wall turbulence (Foggi Rota et al., 2024), nevertheless noticing a significant difference in the dominant oscillation frequency for the turbulence dominated regime. In this case, in facts, the flapping state of the fiber non-trivially relates to the largest scale of the flow and not to the turbulent eddies of comparable size, as found by Olivieri et al. (2021). In the case of a flexible canopy, the individual dynamics of the elements is altered (Fu et al., 2023) and a collective dynamics (honami/monami) emerges on top of that, as measured experimentally by Py et al. (2006) in the case of a crop field driven by the wind. The state-of-the-art direct numerical simulations (DNS) performed by Monti et al. (2023) shed light on the topic, highlighting how such collective motion is decoupled from the structural natural response of the filaments and independent from their rigidity, being driven by the turbulent fluctuations of the flow only. Such causal dependence is coherent with the previous results of the large eddy simulations from Tschisgale et al. (2021) and Wang et al. (2022a).
@@
In this work we investigate the turbulent motion of the fluid and the dynamics of the flexible filaments constituting a dense submerged canopy. We build on top of an extended version of the DNS database generated by Monti et al. (2023), where each filament is modelled individually by means of an immersed boundary method (first introduced by Goldstein et al., 1993) in the Lagrangian formulation described by Yu (2005) and Huang et al. (2007). Consistently with such approach, the filaments are therefore modelled as inextensible beams. We provide a detailed characterisation of the flow above and within the canopy, as well as describing the interaction between the two and comparing them to experimental measurements. We also assess the consequences on the flow of a variation in the filament density and analyse the effects induced by their motion. From a structural standpoint, we characterise the individual motion of the filaments in the canopy and reconcile it with the picture we previously described for single isolated filaments (Foggi Rota et al., 2024). The remaining part of this paper is organised as follows: in §2 we describe the setup of our simulations and the numerical methods employed. The dynamics of the flow is characterised in §3 where, after reporting relevant mean quantities (§3.1) and comparing them to available experimental results (§3.2), we investigate the energy spectra (§3.3) and the Lumley triangle (§3.4) throughout the whole domain. We further explore the interaction between the inner and the outer flow by carrying out a quadrant analysis (§3.5) at the canopy tip. The consequences on the flow of a variation in the density of the filaments are also assessed (§3.6) and the effects of their flapping motion are explored by “freezing” them in their instantaneous deflected configuration (§3.7). In §4 we characterise the structural dynamics of the filaments by analysing their individual motion (§4.1) and describing their flapping state (§4.2). Finally, in §5, we summarise the main outcomes of our investigation and critically discuss their implications, with few remarks on future developments.
@@ 2 Setup and methods
Our simulations are carried out in an open channel described by means of a right handed Cartesian reference frame with the $x$ axis oriented along the streamwise direction and the $y$ axis perpendicular to the bottom wall. The computational domain is therefore a box of volume $L_{x}\times L_{y}\times L_{z}=2\pi H\times H\times 1.5\pi H$, where we enforce the no-slip and no-penetration boundary conditions at the bottom face; no-penetration and stress-free conditions are instead imposed at the top face in the same fashion of Calmet & Magnaudet (2003), while the wall-parallel directions are treated as periodic. The stems of the flexible filaments constituting the canopy are vertically clamped to the bottom wall and protrude upward into the flow as in figure 1. To discretise the fluid flow we adopt an Eulerian grid made of $N_{x}\times N_{y}\times N_{z}=1152\times 384\times 864$ points homogeneously distributed along the periodic directions, while a non homogeneous stretched distribution is adopted along the $y$ axis in order to correctly capture the sharp variation of the velocity at the canopy tip. In particular, we employ a finer and locally uniform resolution in the region containing the canopy, achieving a constant wall-normal spacing $\Delta y/H=0.002$ for $y/H\in[0.0,0.3]$, and smoothly transition to a wider wall-normal spacing above that, attaining $\Delta y/H=0.004$ at $y/H=1$.
@@
[…]
The filaments are represented as mono-dimensional entities, discretised into a line of Lagrangian points, that obey a generalisation of the Euler-Bernoulli beam model allowing for finite deflections, but retaining the inextensibility constraint. In the rigid canopy case (i.e., $Ca=0$) each filament is made of $n_{L}=81$ Lagrangian points, attaining a spatial resolution $\Delta s=h/(n_{L}-1)$ comparable to the Eulerian grid spacing in the wall-normal direction. In the flexible canopy cases, such spatial resolution would impose a too strict constraint on the time step and we therefore reduce $n_{L}$ to 32: this proves acceptable as the flexibility makes the filaments more compliant to the flow and the velocity difference between the two phases is therefore reduced. Different discretisations were tested by Monti et al. (2023) and Foggi Rota et al. (2024), without significant variations in the filament dynamics for the parameters considered here. We use the same approach by Banaei et al. (2020) to model the dynamics of flexible and inextensible filaments. It consists of an extended version of the distributed-Lagrange-multiplier/fictitious-domain (DLM/FD) formulation of the continuum equations introduced by Yu (2005). Denoting with $\mathbf{X}(s,t)$ the position of a point on the neutral axis of a filament as a function of the curvilinear abscissa $s$ and time $t$, and introducing the linear density difference between the filament and the fluid $\Delta\tilde{\rho}=(\rho_{s}-\rho_{f})\pi d^{2}/4$, its structural dynamics is described by
@@
where $T$ is the tension enforcing the inextensibility and $\mathbf{F}$ is the force acting on the filaments computed by the Lagrangian IBM to couple them with the fluid, as described later. We complement equations (3,4) with an appropriate set of boundary conditions, imposing $\mathbf{X}\rvert_{s=0}=\mathbf{X_{0}}$ along with ${\partial\mathbf{X}}/{\partial s}\rvert_{s=0}=(0,1,0)$ at the clamp and ${\partial^{3}\mathbf{X}}/{\partial s^{3}}\rvert_{s=h}={\partial^{2}\mathbf{X}}/{\partial s^{2}}\rvert_{s=h}=\mathbf{0}$ along with $T\rvert_{s=h}=0$ at the free end, and solve them following the approach of Huang et al. (2007). Nevertheless, here, the bending term is treated implicitly as in Banaei et al. (2020) to allow for a larger time step. The set of Lagrangian equations introduced above, in the absence of any external forcing, is passible of a normal mode analysis yielding the natural frequency $f_{nat}=({\beta_{1}}/{(2\pi h^{2})})\sqrt{{\gamma}/{\tilde{\rho_{s}}}}$, related to the natural pulsation $\omega_{1}$ by $f_{nat}={\omega_{1}}/{(2\pi)}$. $\tilde{\rho_{s}}$ is the filament density per unit length, while $\beta_{1}$ is a coefficient approximately equal to $3.516$, determined through the analysis. Writing $\tilde{\rho_{s}}=\rho_{s}\pi d^{2}/4$, where $\rho_{s}$ is the filament density per unit volume, there follows $f_{nat}\approx({3.516}/{(dh^{2})})\sqrt{{\gamma}/{(\rho_{s}\pi^{3})}}$. $f_{nat}$, as noticed by Foggi Rota et al. (2024), plays a significant role in determining the dynamical response of the filaments to the fluid. As in most cases the filaments are flexible and swaying in the flow, they might collide with the wall and with other filaments. We have thus implemented filament-to-filament and filament-to-wall collision models to prevent the stems from crossing each other or the wall while deforming (Snook et al., 2012). Nevertheless, after the extensive testing of different collision models and of their calibration parameters conducted in previous investigations (Monti et al., 2023), the influence of the filament-to-filament collision term on both the filament and the fluid dynamics was found to be very weak, whereas the filament-to-wall interaction model turned out to be necessary only to correctly describe the dynamics of the most flexible filaments, at large values of $Ca$. We thus resort to an inelastic collision model, applying a repulsive force to all the filament points approaching the wall within a range of four grid points.
[…]
The ability of our numerical setup and methods to correctly describe the dynamics of a whole submerged canopy made of flexible filaments, without introducing spurious effects due to the finite size of the domain or the grid, is extensively assessed in Monti et al. (2023). The length of the domain along the homogeneous directions is sufficient to contain the largest turbulent flow structures (Bailey & Stoll, 2013). The vertical size of the domain, instead, is most likely affecting the results. Our simulations, in facts, aim at investigating a submerged canopy rather than a canopy exposed to a boundary layer, for which a significantly higher domain would be needed. In the numerical simulation of submerged canopies, instead, it is customary not to simulate the fluid interface far above their tip, but rather to approximate it as a free slip surface (Sharma & García-Mayoral, 2018; Tschisgale et al., 2021; Wang et al., 2022a; He et al., 2022; Löhrer & Fröhlich, 2023); its distance from the bottom wall thus becomes a parameter of the simulations. In this work, we follow such practice.
Canopy flows have been investigated experimentally since the origins of the field, and a broad range of measurements is now available (e.g., Gao et al., 1989; Okamoto & Nezu, 2009; Nicolai et al., 2020). Nevertheless, the comparison to simulation results is often hindered by the inevitable differences between numerical and experimental setups, especially in the canopy arrangement and flow parameters. Experiments are often performed at significantly higher Reynolds numbers than simulations, and they are frequently characterised by the presence of secondary flows. Despite these difficulties, we found that our setup compares well to that adopted by Shimizu et al. (1992), where a rigid canopy of height $h=0.65H$ and solidity $\lambda=0.41$ is exposed to a turbulent channel flow at $Re_{b}=7070$. Therefore, after purposely simulating the flow within and over a rigid canopy matching the same parameters (further details in the supplementary information of Monti et al., 2023), we contrast the computed mean flow profile and Reynolds shear stress with their measurements in figure 2. This comparison confirms the ability of our code to correctly describe the back-reaction of the structure on the fluid.
Figure 2: Mean streamwise velocity profile (panel $a$) and Reynolds shear stress (panel $b$) in and above a rigid canopy with $h=0.65H$ and $\lambda=0.41$, at $Re_{b}=7070$. Red stars denote the experimental measurements of Shimizu et al. (1992), while black lines are the outcome of a direct numerical simulation matching the experimental parameters, performed with our code.